Suppose that a firms short-run production function is F(K,L) = L+VKL, where capital is fixed at K = 4. The rental rate of caIs my answer correct?
Answer #1

F=L+ UKL k = 4 wage rate (w) = 24 r = rental rate = 24 & F = 48 48= L+S4L = L+ FL L+250-48 =0 L+85-65 -48 =0 [SL r A - 6(JL +